OpenStudy (anonymous):

Solve the system of equations and enter the solution as an ordered pair. 2y = 4x - 16 5y = 4x + 8

OpenStudy (anonymous):

Subtract the equations. -3y = -24 y = 8 2(8) = 4x - 16 16 = 4x - 16 32 = 4x x = 8 (8, 8)